# Autodeleveraging: Impossibilities and Optimization

### v1

@@ B.6 Randomized constructions for moral-hazard examples
Appendix C Theoretical Properties of Capped Pro-Rata
We formalize the theoretical properties of the capped pro-rata rule. We note that the capped pro-rata algorithm in Algorithm 1 is a standard water-filling algorithm [BoydVandenberghe2004]. The most similar known prior work to this appendix is the study of how such algorithms provide sybil resistance in concave games in decentralized systems [johnson2023concave]. One can view our result as a generalization of this result.
@@ Appendix C Theoretical Properties of Capped Pro-Rata
Fix time $T$, state $\mathcal{P}_{n}$, winners $\mathcal{W}_{T}$, and equities $e_{i}=(e_{T}(\mathfrak{p}_{i}))_{+}$. Let $s_{\pi,i}=(1-h_{\pi,i})e_{i}$ be the surviving equity of position $i$ under ADL policy $\pi$. We say that a feasible ADL rule $\pi$2.4) satisfies:
@@
    Sybil resistance: Outcomes are invariant to splitting/merging accounts. For any split $e_{i}=\sum_{a=1}^{r}z_{a}$, the sum of survivors equals the original survivor: $\sum_{a}s^{(i\to z)}_{\pi,a}=s_{\pi,i}$.
@@
    Scale invariance: $s_{\pi}(ce;\theta_{\pi})=c\,s_{\pi}(e;\theta_{\pi})$ for $c>0$.
@@
    Monotonicity: If $e_{1}\geq\dots\geq e_{k}$, then $s_{\pi,1}\geq\dots\geq s_{\pi,k}$.
@@
    Interior regularity: The map $(e,H)\mapsto s_{\pi}(e;H)$ is $C^{1}$ on the interior, i.e., for $e_{i}>0$ for all $i$ and $0<H<\sum_{i}e_{i}$.
@@
Proposition C.1 (Uniqueness of the Pro-Rata Rule).
If a feasible ADL policy $\pi$ satisfies the fairness properties, then $s_{\pi}(e;H)=s^{\mathrm{PR}}(e;H)$ for all feasible inputs.
@@
Fix a feasible budget $H$ (e.g. maximum value of $\theta_{n}D_{T}$) and write $\beta_{i}\in(0,1]$ for the haircut cap on winner $i$ (as in the capped pro-rata rule (29)), and sort so that $e_{1}\geq\cdots\geq e_{k}$. View the haircuts as a function of the realized budget $b\in[0,H]$ and write $h_{i}(b)$ for the haircut on winner $i$ when total budget $b$ has been allocated. For each $b$, define the active set
@@
of winners whose caps are not yet binding. Since there are only finitely many caps $\beta_{i}$, there exists a partition $0=b_{0}<b_{1}<\cdots<b_{L}=H$ such that on each open interval $I_{\ell}:=(b_{\ell-1},b_{\ell})$ the active set is constant. Fix one such interval $I=(b_{-},b_{+})$ and write $A=A(b)$ for any $b\in I$. On this interval, by scale invariance on $A$ (per–unit budget increases all active haircuts at the same rate) and feasibility,
$\frac{dh_{i}}{db}\;=\;\frac{1}{\sum_{j\in A}e_{j}}\quad(i\in A),\qquad\frac{dh_{i}}{db}=0\quad(i\notin A).$
@@
$\tilde{h}_{i}(b)\;=\;\begin{cases}h_{i}(b_{-})\;+\;\dfrac{b-b_{-}}{\sum_{j\in A}e_{j}}&(i\in A),\\[4.0pt] h_{i}(b_{-})&(i\notin A).\end{cases}$
@@
$b\;=\;\sum_{i\notin A}e_{i}h_{i}(b_{-})\;+\;\sum_{i\in A}e_{i}\min\{\eta_{I}(b),\beta_{i}\},$
which, for fixed $b\in I$, is a continuous strictly increasing function of $\eta_{I}(b)$ as long as $A\neq\emptyset$. Thus for each $b\in I$ there is a unique $\eta_{I}(b)$ solving the budget identity. In particular, on the interior interval where no caps bind ($A=\{1,\dots,k\}$ and $b\in(0,W_{T})$), we have $h_{i}(b)=\eta_{I}(b)$ for all $i$ so the budget identity reduces to
$b\;=\;\sum_{i=1}^{k}e_{i}h_{i}(b)\;=\;\eta_{I}(b)\sum_{i=1}^{k}e_{i}\;=\;\eta_{I}(b)\,W_{T},$
which implies $\eta_{I}(b)=b/W_{T}$. Thus on this interval $s_{\pi,i}=(1-\eta_{I}(b))e_{i}=(1-b/W_{T})e_{i}$, i.e., proportional to equity.
Sybil resistance implies allocations depend only on total equity: splitting $e_{i}=\sum_{a}z_{a}$ leaves $\sum_{a}z_{a}h_{a}(b)$ and hence the survivor $\sum_{a}(1-h_{a}(b))z_{a}$ unchanged for each $b$, so the proportional form on interior intervals is preserved under arbitrary splits. Monotonicity further restricts how indices can exit the active set as $b$ increases. At an endpoint $b_{\ell}$ where $\eta_{I}(b_{\ell})$ first hits some cap $\beta_{m}$, all indices $j\geq m$ with $\beta_{j}\leq\beta_{m}$ must saturate together; otherwise we would have $e_{m}\geq e_{j}$ but $(1-h_{m}(b_{\ell}))e_{m}<(1-h_{j}(b_{\ell}))e_{j}$, violating monotonicity. Thus, as we pass from $I_{\ell}$ to $I_{\ell+1}$, a (possibly empty) tail $\{j>m\}$ leaves $A$, contributing a fixed amount $\sum_{j>m}e_{j}\beta_{j}$ to the budget, and the same water-filling argument applies on the remaining head with reduced budget.
Concatenating the solution across all intervals $I_{\ell}$ yields the reverse–waterfilling form $h_{i}=\min\{\eta^{\star},\beta_{i}\}$ with $\eta^{\star}$ chosen so that $\sum_{i}e_{i}h_{i}=H$, which is exactly capped pro–rata. Uniqueness follows either from the strict convexity of the Euclidean projection onto $\{h\in[0,1]^{k}:\ \sum_{i}e_{i}h_{i}=H\}$ or from the monotone one–dimensional search that defines $\eta^{\star}$. ∎
@@
We now formalize the convex-welfare interpretation of capped pro-rata from §6. Fix time $T$, winners $\mathcal{W}_{T}$ with equities $e_{i}=e_{T}(\mathfrak{p}_{i})_{+}$, and effective caps $\beta_{i}=\min\{\overline{h}_{i},1-\underline{e}_{i}/e_{T}(\mathfrak{p}_{i})\}$ defined by the haircut and equity constraints (22)–(23). Let $B_{T}=\theta_{\pi}D_{T}(\mathcal{P}_{n})$ denote the haircut budget from (30), and let $\phi:[0,1]\to{\mbox{\bf R}}$ be a strictly convex increasing function representing per-unit haircut disutility as in (31). We consider choosing haircuts $h=(h_{i})_{i\in\mathcal{W}_{T}}$ to minimize the equity-weighted total disutility $\sum_{i\in\mathcal{W}_{T}}e_{i}\phi(h_{i})$ subject to the per-account bounds $0\leq h_{i}\leq\beta_{i}$ and the aggregate budget constraint $\sum_{i\in\mathcal{W}_{T}}e_{i}h_{i}=B_{T}$.
Proposition C.2 (Convex optimality).
@@
$\min_{h}\ \sum_{i\in\mathcal{W}_{T}}e_{i}\,\phi(h_{i})\quad\text{s.t.}\quad\sum_{i\in\mathcal{W}_{T}}e_{i}h_{i}=B_{T},\quad 0\leq h_{i}\leq\beta_{i}$
is the capped pro-rata rule (29), i.e., $h_{\pi_{CP},i}^{\star}=\min\{\eta^{\star},\beta_{i}\}$, where $\eta^{\star}$ is chosen so that $\sum_{i\in\mathcal{W}_{T}}e_{i}h_{\pi_{CP},i}^{\star}=B_{T}$.
@@
The optimization problem is convex with a strictly convex objective and linear constraints, so any point satisfying the Karush–Kuhn–Tucker (KKT) conditions is the unique global minimizer [BoydVandenberghe2004, Ch. 5]. For $B_{T}$ in the interior of the feasible region ($0<B_{T}<C(\beta)$, where $C(\beta)$ is defined in (28)), Slater’s condition holds, so KKT conditions are necessary and sufficient. The Lagrangian is
$\mathcal{L}(h,\lambda,\mu,\nu)\;=\;\sum_{i\in\mathcal{W}_{T}}e_{i}\phi(h_{i})+\lambda\!\left(\sum_{i\in\mathcal{W}_{T}}e_{i}h_{i}-B_{T}\right)+\sum_{i\in\mathcal{W}_{T}}\mu_{i}(h_{i}-\beta_{i})-\sum_{i\in\mathcal{W}_{T}}\nu_{i}h_{i},$
@@
$e_{i}\phi^{\prime}(h_{i})+\lambda e_{i}+\mu_{i}-\nu_{i}=0\quad(i\in\mathcal{W}_{T}),$
@@
The budget constraint $\sum_{i\in\mathcal{W}_{T}}e_{i}h_{i}=B_{T}$ then reduces to finding $c$ such that
$\sum_{i\in\mathcal{W}_{T}}e_{i}\min\{c,\beta_{i}\}=B_{T}.$
The left-hand side is continuous and strictly increasing in $c$ on $[0,1]$ as long as some $\beta_{i}>0$, so there is a unique $c=\eta^{\star}$ solving this equation. Thus every KKT point has the capped pro-rata form $h_{i}=\min\{\eta^{\star},\beta_{i}\}$ with $\eta^{\star}$ determined by the budget, which is exactly the rule in (29). Strict convexity of the objective implies this KKT point is the unique global minimizer, proving the claim (see also the standard water-filling derivations in [BoydVandenberghe2004, Ch. 5]). ∎
Appendix D Algorithms for Pro-Rata Haircut Rules
This appendix collects the explicit water-filling procedures for both the capped pro-rata rule from Section˜6 and the Risk-Aware Pro-Rata (RAP) rule from Section˜7. Each algorithm takes the winner equities $\{e_{i}\}$, effective caps $\{\beta_{i}\}$, and the target budget $B_{T}=\theta_{\pi}D_{T}(\mathcal{P}_{n})$, returning the optimal haircut vector $h$ or declaring infeasibility if the aggregate capacity $\sum_{i}e_{i}\beta_{i}$ is insufficient.
D.1 Capped Pro-Rata Water-Filling
The procedure below enforces the constraints in Equations˜29 and 30 by leveling caps until the target budget is met.

### v3

@@ C.6 Randomized constructions for moral-hazard examples
Appendix D Theoretical Properties of Capped Pro-Rata
We formalize the theoretical properties of the capped pro-rata rule. We note that the capped pro-rata algorithm in Algorithm 1 is a standard water-filling algorithm [29]. The most similar known prior work to this appendix is the study of how such algorithms provide sybil resistance in concave games in decentralized systems [74]. One can view our result as a generalization of this result.
@@ Appendix D Theoretical Properties of Capped Pro-Rata
Fix time $T$, state $\mathcal{P}_{n}$, winners $\mathcal{W}_{T}$, and endowments $w_{i}=w_{T,i}$ (under PNL-only, $w_{i}=(\mathrm{PNL}_{T,i})_{+}$). Let $s_{\pi,i}=(1-h_{\pi,i})w_{i}$ be the surviving endowment of position $i$ under ADL policy $\pi$. The post-ADL equity is $e^{\prime}_{\pi,i}=c_{i,T}+\mathrm{PNL}_{i,T}-h_{\pi,i}w_{i}=c_{i,T}+(1-h_{\pi,i})w_{i}+(\mathrm{PNL}_{i,T})_{-}$ (where $(\mathrm{PNL}_{i,T})_{-}$ is the negative PNL component, if any). We say that a feasible ADL rule $\pi$2.4) satisfies:
@@
    Sybil resistance: Outcomes are invariant to splitting/merging accounts. For any split $w_{i}=\sum_{a=1}^{r}z_{a}$, the sum of survivors equals the original survivor: $\sum_{a}s^{(i\to z)}_{\pi,a}=s_{\pi,i}$.
@@
    Scale invariance: $s_{\pi}(cw;\theta_{\pi})=c\,s_{\pi}(w;\theta_{\pi})$ for $c>0$.
@@
    Monotonicity: If $w_{1}\geq\dots\geq w_{k}$, then $s_{\pi,1}\geq\dots\geq s_{\pi,k}$ (and consequently, if cash components are comparable, post-ADL equities preserve ordering).
@@
    Interior regularity: The map $(w,H)\mapsto s_{\pi}(w;H)$ is $C^{1}$ on the interior, i.e., for $w_{i}>0$ for all $i$ and $0<H<\sum_{i}w_{i}$.
@@
Queue score dependence: Sybil resistance vs monotonicity.
The queue rule’s Sybil behavior depends on the ranking score, while monotonicity can fail even when Sybil resistance holds.
Proposition D.1 (Queue Sybil resistance for score-preserving splits).
Consider the wealth-space queue rule $\pi_{Q}$ with budget $B$, winner endowments $w$, and ranking scores $r_{i}$. Fix winner $i$ with a unique score level ($r_{i}\neq r_{j}$ for all $j\neq i$). Suppose $i$ is split into subaccounts with endowments $z_{1},\dots,z_{m}>0$, $\sum_{a=1}^{m}z_{a}=w_{i}$, and each child keeps the parent score ($r_{i,a}=r_{i}$). Write $x_{\pi_{Q},i}(B):=w_{i}-s_{\pi_{Q},i}(B)$ for seizure at budget $B$, and $\tilde{x}_{\pi_{Q},a}(B)$ for seizure on split child $a$. Then aggregate seizure is invariant:
$\sum_{a=1}^{m}\tilde{x}_{\pi_{Q},a}(B)=x_{\pi_{Q},i}(B).$
Hence queue is Sybil resistant for such splits.
@@
Let $H_{i}=\{j:\ r_{j}>r_{i}\}$ and define the residual budget when the queue reaches score level $r_{i}$:
$R_{i}=\Big(B-\sum_{j\in H_{i}}w_{j}\Big)_{+}.$
Because $r_{i}$ is unique, no other account is processed between $H_{i}$ and account $i$. In the unsplit instance, queue seizure is
$x_{\pi_{Q},i}(B)=\min\{w_{i},\ R_{i}\}.$
In the split instance, the same higher-score set $H_{i}$ is processed first, so the residual is still $R_{i}$ when the split children are reached. The queue then seizes greedily across the children, whose total endowment is $\sum_{a}z_{a}=w_{i}$, hence
$\sum_{a=1}^{m}\tilde{x}_{\pi_{Q},a}(B)=\min\Big\{\sum_{a=1}^{m}z_{a},\ R_{i}\Big\}=\min\{w_{i},\ R_{i}\}=x_{\pi_{Q},i}(B).$
Corollary D.2 (Hyperliquid score is Sybil resistant under proportional splits).
Hyperliquid’s documented ADL score [72],
$r=\frac{\mathrm{mark}}{\mathrm{entry}}\cdot\frac{\mathrm{notional}}{\mathrm{account\ value}},$
is homogeneous of degree zero in $(\mathrm{notional},\mathrm{account\ value})$ under proportional scaling: for any $\alpha>0$,
$r(\alpha\,\mathrm{notional},\alpha\,\mathrm{account\ value})=r(\mathrm{notional},\mathrm{account\ value}).$
Therefore proportional account splitting preserves score and, under Proposition D.1’s no-tie assumption at the split score level, preserves aggregate queue seizure (Sybil resistance). See also the worked example in [1].
Proof.
The ratio $\mathrm{notional}/\mathrm{account\ value}$ is unchanged by common scaling, and the mark/entry term is unchanged by splitting, so each split child has the same score as the parent. Apply Proposition D.1. ∎
Proposition D.3 (Queue monotonicity failure).
There exist two winners with $w_{1}>w_{2}>0$, queue scores $r_{1}>r_{2}$, and budget $B\in(w_{1}-w_{2},w_{1}]$ such that queue violates monotonicity:
$s^{\prime}_{1}<s^{\prime}_{2},$
where $s^{\prime}_{i}=w_{i}-x_{i}$ is post-ADL surviving endowment.
Proof.
With $r_{1}>r_{2}$, queue processes winner $1$ first. For $B\leq w_{1}$, queue seizes $x_{1}=B$ and $x_{2}=0$. Hence
$s^{\prime}_{1}=w_{1}-B,\qquad s^{\prime}_{2}=w_{2}.$
If $B>w_{1}-w_{2}$, then $w_{1}-B<w_{2}$, so $s^{\prime}_{1}<s^{\prime}_{2}$ despite $w_{1}>w_{2}$ pre-ADL. ∎
Absolute-score vulnerability.
If the score is an absolute quantity (e.g. $r_{i}=\mathrm{PNL}_{i}$), splitting changes child scores and can move a position deeper in the queue. This is the standard Sybil-vulnerable queue setting.
Proposition D.4 (Uniqueness of the Pro-Rata Rule).
If a feasible ADL policy $\pi$ satisfies the fairness properties, then $s_{\pi}(w;H)=s^{\mathrm{PR}}(w;H)$ for all feasible inputs, where $s$ is the surviving endowment.
Proof.
Fix a feasible budget $H$ (e.g. maximum value of $\theta_{n}D_{T}$) and write $\beta_{i}\in(0,1]$ for the haircut cap on winner $i$ (as in the capped pro-rata rule (33)), and sort so that $w_{1}\geq\cdots\geq w_{k}$. View the haircuts as a function of the realized budget $b\in[0,H]$ and write $h_{i}(b)$ for the haircut on winner $i$ when total budget $b$ has been allocated. For each $b$, define the active set
@@
of winners whose caps are not yet binding. Since there are only finitely many caps $\beta_{i}$, there exists a partition $0=b_{0}<b_{1}<\cdots<b_{L}=H$ such that on each open interval $I_{\ell}:=(b_{\ell-1},b_{\ell})$ the active set is constant. Fix one such interval $I=(b_{-},b_{+})$ and write $A=A(b)$ for any $b\in I$. On this interval, by scale invariance on $A$ (per–unit budget increases all active haircuts at the same rate) and feasibility (budget constraint in endowment space),
$\frac{dh_{i}}{db}\;=\;\frac{1}{\sum_{j\in A}w_{j}}\quad(i\in A),\qquad\frac{dh_{i}}{db}=0\quad(i\notin A).$
@@
$\tilde{h}_{i}(b)\;=\;\begin{cases}h_{i}(b_{-})\;+\;\dfrac{b-b_{-}}{\sum_{j\in A}w_{j}}&(i\in A),\\[4.0pt] h_{i}(b_{-})&(i\notin A).\end{cases}$
@@
$b\;=\;\sum_{i\notin A}w_{i}h_{i}(b_{-})\;+\;\sum_{i\in A}w_{i}\min\{\eta_{I}(b),\beta_{i}\},$
which, for fixed $b\in I$, is a continuous strictly increasing function of $\eta_{I}(b)$ as long as $A\neq\emptyset$. Thus for each $b\in I$ there is a unique $\eta_{I}(b)$ solving the budget identity. In particular, on the interior interval where no caps bind ($A=\{1,\dots,k\}$ and $b\in(0,U_{T})$ where $U_{T}=\sum_{i}w_{i}$), we have $h_{i}(b)=\eta_{I}(b)$ for all $i$ so the budget identity reduces to
$b\;=\;\sum_{i=1}^{k}w_{i}h_{i}(b)\;=\;\eta_{I}(b)\sum_{i=1}^{k}w_{i}\;=\;\eta_{I}(b)\,U_{T},$
which implies $\eta_{I}(b)=b/U_{T}$. Thus on this interval $s_{\pi,i}=(1-\eta_{I}(b))w_{i}=(1-b/U_{T})w_{i}$, i.e., proportional to endowment.
Sybil resistance implies allocations depend only on total endowment: splitting $w_{i}=\sum_{a}z_{a}$ leaves $\sum_{a}z_{a}h_{a}(b)$ and hence the survivor $\sum_{a}(1-h_{a}(b))z_{a}$ unchanged for each $b$, so the proportional form on interior intervals is preserved under arbitrary splits. Monotonicity further restricts how indices can exit the active set as $b$ increases. At an endpoint $b_{\ell}$ where $\eta_{I}(b_{\ell})$ first hits some cap $\beta_{m}$, all indices $j\geq m$ with $\beta_{j}\leq\beta_{m}$ must saturate together; otherwise we would have $w_{m}\geq w_{j}$ but $(1-h_{m}(b_{\ell}))w_{m}<(1-h_{j}(b_{\ell}))w_{j}$, violating monotonicity. Thus, as we pass from $I_{\ell}$ to $I_{\ell+1}$, a (possibly empty) tail $\{j>m\}$ leaves $A$, contributing a fixed amount $\sum_{j>m}w_{j}\beta_{j}$ to the budget, and the same water-filling argument applies on the remaining head with reduced budget.
Concatenating the solution across all intervals $I_{\ell}$ yields the reverse–waterfilling form $h_{i}=\min\{\eta^{\star},\beta_{i}\}$ with $\eta^{\star}$ chosen so that $\sum_{i}w_{i}h_{i}=H$, which is exactly capped pro–rata. Uniqueness follows either from the strict convexity of the Euclidean projection onto $\{h\in[0,1]^{k}:\ \sum_{i}w_{i}h_{i}=H\}$ or from the monotone one–dimensional search that defines $\eta^{\star}$. ∎
@@
We now formalize the convex-welfare interpretation of capped pro-rata from §6. Fix time $T$, winners $\mathcal{W}_{T}$ with endowments $w_{i}=w_{T,i}$ (under PNL-only, $w_{i}=(\mathrm{PNL}_{T,i})_{+}$), and effective caps $\beta_{i}=\min\{\overline{h}_{i},1-(\underline{e}_{i}-c_{i,T})/w_{i,T}\}$ defined by the haircut and equity constraints (26)–(27), where the equity constraint is reinterpreted in terms of minimum post-ADL endowment. Let $B_{T}=\theta_{\pi}D_{T}(\mathcal{P}_{n})$ denote the haircut budget from (34), and let $\phi:[0,1]\to{\mbox{\bf R}}$ be a strictly convex increasing function representing per-unit haircut disutility as in (35). We consider choosing haircuts $h=(h_{i})_{i\in\mathcal{W}_{T}}$ to minimize the endowment-weighted total disutility $\sum_{i\in\mathcal{W}_{T}}w_{i}\phi(h_{i})$ subject to the per-account bounds $0\leq h_{i}\leq\beta_{i}$ and the aggregate budget constraint $\sum_{i\in\mathcal{W}_{T}}w_{i}h_{i}=B_{T}$ (budget constraint in endowment space).
Proposition D.5 (Convex optimality).
@@
$\min_{h}\ \sum_{i\in\mathcal{W}_{T}}w_{i}\,\phi(h_{i})\quad\text{s.t.}\quad\sum_{i\in\mathcal{W}_{T}}w_{i}h_{i}=B_{T},\quad 0\leq h_{i}\leq\beta_{i}$
is the capped pro-rata rule (33), i.e.,
$h_{\pi_{CP},i}^{\star}=\min\{\eta^{\star},\beta_{i}\},$
where $\eta^{\star}$ satisfies $\sum_{i\in\mathcal{W}_{T}}w_{i}h_{\pi_{CP},i}^{\star}=B_{T}$.
@@
The optimization problem is convex with a strictly convex objective and linear constraints, so any point satisfying the Karush–Kuhn–Tucker (KKT) conditions is the unique global minimizer [29, Ch. 5]. For $B_{T}$ in the interior of the feasible region ($0<B_{T}<C(\beta)$, where $C(\beta)=\sum_{i}w_{i}\beta_{i}$ is defined in (32)), Slater’s condition holds, so KKT conditions are necessary and sufficient. The Lagrangian is
$\mathcal{L}(h,\lambda,\mu,\nu)\;=\;\sum_{i\in\mathcal{W}_{T}}w_{i}\phi(h_{i})+\lambda\!\left(\sum_{i\in\mathcal{W}_{T}}w_{i}h_{i}-B_{T}\right)+\sum_{i\in\mathcal{W}_{T}}\mu_{i}(h_{i}-\beta_{i})-\sum_{i\in\mathcal{W}_{T}}\nu_{i}h_{i},$
@@
$w_{i}\phi^{\prime}(h_{i})+\lambda w_{i}+\mu_{i}-\nu_{i}=0\quad(i\in\mathcal{W}_{T}),$
@@
The budget constraint $\sum_{i\in\mathcal{W}_{T}}w_{i}h_{i}=B_{T}$ then reduces to finding $c$ such that
$\sum_{i\in\mathcal{W}_{T}}w_{i}\min\{c,\beta_{i}\}=B_{T}.$
The left-hand side is continuous and strictly increasing in $c$ on $[0,1]$ as long as some $\beta_{i}>0$, so there is a unique $c=\eta^{\star}$ solving this equation. Thus every KKT point has the capped pro-rata form $h_{i}=\min\{\eta^{\star},\beta_{i}\}$ with $\eta^{\star}$ determined by the budget, which is exactly the rule in (33). Strict convexity of the objective implies this KKT point is the unique global minimizer, proving the claim (see also the standard water-filling derivations in [29, Ch. 5]). ∎
Appendix E Algorithms for Pro-Rata Haircut Rules
This appendix collects the explicit water-filling procedures for both the capped pro-rata rule from Section˜6 and the Risk-Aware Pro-Rata (RAP) rule from Section˜7. Each algorithm takes the winner endowments $\{w_{i}\}$ (under PNL-only, positive PNL), effective caps $\{\beta_{i}\}$, and the target budget $B_{T}=\theta_{\pi}D_{T}(\mathcal{P}_{n})$, returning the optimal haircut vector $h$ or declaring infeasibility if the aggregate capacity $\sum_{i}w_{i}\beta_{i}$ is insufficient.
E.1 Capped Pro-Rata Water-Filling
The procedure below enforces the constraints in Equations˜33 and 34 by leveling caps until the target budget is met.